OpenStudy (anonymous):

@study100 Help me on these two questions please

OpenStudy (anonymous):

I'll try ;u;

OpenStudy (anonymous):

I did what we did in the previous ones right and took "y" out of the equation and took the integral based on x

OpenStudy (anonymous):

\[\large \large \frac{ dy }{ \sqrt{y} } = \frac{ dx }{ \sqrt{x}}\]

OpenStudy (anonymous):

and just intergrate both sides : DDD

OpenStudy (anonymous):

\[\frac{ 2 }{\sqrt{\frac{1}{x}}}+ C\]

OpenStudy (anonymous):

\[\large 2y^{1/2} = 2 x ^{1/2} + C\]

OpenStudy (anonymous):

the other side is \[\frac{ y^2}{2} +C\]

OpenStudy (anonymous):

lol im wrong then

OpenStudy (anonymous):

f(x) =\[2\sqrt{x}\]

OpenStudy (anonymous):

well integrate we have \[\large 2 y ^{-1/2 +1}\] = \[2 x ^ {-1/2 + 1} + C???\]

OpenStudy (anonymous):

simplify I got \[\sqrt{y} = \sqrt{x} +C\]

OpenStudy (anonymous):

nice i simplified and got the same thing

OpenStudy (anonymous):

one more question and thats it

OpenStudy (anonymous):

Ok! xD I almost have to eat dinner too! \[\large \frac{ dy }{ y } = 3x dx\]

OpenStudy (anonymous):

\[\large \ln (y) =\frac{ 3x^{2} }{ 2 } + C\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

should I exponentiate it? o.o like e^lny to get y alone?

OpenStudy (anonymous):

\[\large e^{lny} = e^{3x^{2}/2+C}\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

\[y = Ce^{3x^{2}/2}\]

OpenStudy (anonymous):

the answer is